H2O2(aq)⟶H2O(l) + 1/2 O2(g)

a. The above chemical equation is for a first-order reaction. At 302 K , the rate constant equals 8.5×10−4s−1. Calculate the half-life at this temperature. Express your answer to two significant figures and include the appropriate units.

b. If the activation energy for this reaction is 75 kJ/mol, at what temperature would the reaction rate be doubled? Express your answer as an integer and include the appropriate units.
Answer #1

a)

half life = 0.693 /rate constant

            = 0.693 / 8.5 x 10^-4

           = 815.3 sec

half life = 8.2 x 10^2 sec

b)

ln (k2 / k1) = Ea / R [1 / T1 - 1 / T2 ]

ln 2 = 75 / 8.314 x 10^-3 [1 / 302 - 1 / T2 ]

7.68 x 10^-5 = 3.31 x 10^-3 - 1/ T2

T2 = 309 K

temperature = 309 K